r =1.5 + 0.75 π

1. Using the formula from above, what is the Aggregate Demand function?

2-Suppose real money demand is describe by equation L(i,Y) = 8 - 2i + 0.5Y and assume that income is Y = $10 billion and real money supply is $7 Billion. What is the money demand formula?

3-What is the equilibrium (nominal interest rate?

4-If the central bank increases the real money supply to $4 billon calculate the new interest rate?
1
Expert's answer
2014-11-14T14:52:41-0500
1. Aggregate demand (AD) is the total demand for final goods and services in an economy at a given time.
AD = C + I + G + NX
2. Real money demand L(i,Y) = 8 - 2i + 0.5Y, Y = $10 billion, real money supply MS = $7 Billion.
L = MS, so:
8 - 2i + 0.5*10 = 7
The money demand formula is L = 8 - 2i + 5
3-What is the equilibrium (nominal interest rate?
L = MS, so:
8 - 2i + 0.5*10 = 7
2i = 6,
i = 3%
4. If the central bank increases the real money supply to $4, the new interest rate will be:
8 - 2i + 5 = 4
2i = 9
i = 4.5%